# Accurate Models of NVIDIA Tensor Cores

### v1

@@ V-B Multi-Word Algorithms for Emulating High-Precision GEMM on Tensor Cores
Interestingly, for single-word arithmetic, all models match up except when the problem size increases close to $n=10^{6}$. For double- and triple-word arithmetic, the results show that the V100 tensor core provides two orders of magnitude lower error at $n=10^{6}$. This may be caused by B200 tensor core model having more accurate accumulator with ${n_{\mathrm{eab}}}=2$, which is rounded to zero, making the error larger than the V100 tensor core’s result. This can occur if the V100 result is above the reference result, whilst the B200 result is closer to the reference, but below it, so that RZ pushes it further below. We have tested this hypothesis by enabling RN in the B200, which demonstrates an improvement. Further analysis of this behavior is out of the scope of this paper and we leave it for future work. Finally, the L40S GPU is, as expected, the least accurate because of the 13 fractional bits in the accumulator as discussed in Section IV-A5.
@@ VI Conclusion
We have discussed the research behind the development of the MATLAB Tensor Core v0.1 toolbox. The toolbox contains various NVIDIA GPU tensor core models, as well as a parameterised model that can be used to instantiate variants of real tensor cores for research purposes. The models were verified against the GPU hardware by large-scale randomised testing and model refinement. The proposed toolbox includes tensor core models for NVIDIA A2, A30, A100, Ada 1000 RTX, L40S, H100, H200, and B200, supporting all input precision formats available in CUDA, except binary64. For binary64, NVIDIA GPU tensor cores behave as sequential chains of IEEE-compliant FMAs; therefore, Matlab’s built-in fma function can be used to emulate such tensor core behaviour. In addition to the fixed models, we provide a custom tensor-core model that enables users to simulate arbitrary configurations by adjusting the FMA size, the number of extra bits used for aligning significands, and the rounding mode (supporting RNE, RZ, RD, and RU). For matrix multiplication, the toolbox utilises the Matlab Parallel Computing Toolbox to create a parallel pool of workers (using the default profile). This dispatches inner products within a GEMM across multiple Matlab engines.

### v2

@@ V-B Multi-Word Algorithms for Emulating High-Precision GEMM on Tensor Cores
Interestingly, for single-word arithmetic, all models match up except when the problem size increases close to $n=10^{6}$. For double- and triple-word arithmetic, the results show that the V100 tensor core provides two orders of magnitude lower error at $n=10^{6}$. This may be caused by B200 tensor core model having more accurate accumulator with ${n_{\mathrm{eab}}}=2$, which is rounded to zero, making the error larger than the V100 tensor core’s result. This can occur if the V100 result is above the reference result, whilst the B200 result is closer to the reference, but below it, so that RZ pushes it further below. We have tested this hypothesis by enabling RN in the B200, which demonstrates an improvement. Further analysis of this behavior is out of the scope of this paper and we leave it for future work. Finally, the L40S GPU is, as expected, the least accurate because of the 13 fractional bits in the accumulator as discussed in Section IV-A4.
@@ VI Conclusion
We have discussed the research behind the development of the MATLAB Tensor Core v0.2 toolbox. The toolbox contains various NVIDIA GPU tensor core models, as well as a parameterised model that can be used to instantiate variants of real tensor cores for research purposes. The models were verified against the GPU hardware by large-scale randomised testing and model refinement. The proposed toolbox includes tensor core models for NVIDIA A2, A30, A100, Ada 1000 RTX, L40S, H100, H200, and B200, supporting all input precision formats available in CUDA, except binary64. For binary64, NVIDIA GPU tensor cores behave as sequential chains of IEEE-compliant FMAs; therefore, Matlab’s built-in fma function can be used to emulate such tensor core behaviour. In addition to the fixed models, we provide a custom tensor-core model that enables users to simulate arbitrary configurations by adjusting the FMA size, the number of extra bits used for aligning significands, and the rounding mode (supporting RNE, RZ, RD, and RU). For matrix multiplication, the toolbox utilises the Matlab Parallel Computing Toolbox to create a parallel pool of workers (using the default profile). This dispatches inner products within a GEMM across multiple Matlab engines.